OpenStudy (anonymous):

If f(x) = |x – 5| + 2, find f(3). 10 6 4 0

OpenStudy (whpalmer4):

\[f(x) = |x-5| + 2\]\[f(3) = |3-5|+2 = |-2|+2 = 2+2 =\]

OpenStudy (anonymous):

so i solve the second equation and that my answer?

OpenStudy (whpalmer4):

Yeah, all you have to do is add 2 and 2 :-)